1. An object whose mass is 0.05 kg is moving at a rate of 100 m/s. What force will be needed to stop the object in 0.14 s?
1
Expert's answer
2021-01-10T18:27:19-0500

"F=\frac pt=\frac{mv}{t}=\frac{0.05\cdot 100}{0.14}≈35.7~N."